OpenStudy (anonymous):

√(2b - a+b)^2 + (2c - c)^2

terenzreignz (terenzreignz):

@LoveYou*69

terenzreignz (terenzreignz):

Oh, there you are...

terenzreignz (terenzreignz):

What are we supposed to do, simplify?

terenzreignz (terenzreignz):

\[\Large \sqrt{(2b - a+b)^2 + (2c -c)^2}\]

OpenStudy (anonymous):

I'm supposed to simplify that and √(2a - a+b)^2 + (0 - c)^2 enough that I can tell they are equal. I can do this one if u help me with that one first :P

terenzreignz (terenzreignz):

Of course :) 2b - a + b is just 3b - a, correct?

OpenStudy (anonymous):

Yeah, I think, and 2c -c is c so is it √(3b-a)^2 + (c)^2 ??

terenzreignz (terenzreignz):

That is correct... \[\Large = \sqrt{(3b-a)^2 +(c)^2}\]

OpenStudy (anonymous):

Okay thank you.

terenzreignz (terenzreignz):

Now, what about the other one?

OpenStudy (anonymous):

I can do the other one lol hold on

terenzreignz (terenzreignz):

\[\Large \sqrt{(2b - a+b)^2 + (0 -c)^2}\]

OpenStudy (anonymous):

MU = √(a + b)^2 + (-c)^2 MU = √(a + b)^2 + c ?

terenzreignz (terenzreignz):

Hang on, this one... \[\Large \sqrt{(2b - a+b)^2 + (2c -c)^2}\] simplifies into \[\Large \sqrt{(3b-a)^2 + (c)^2}\] yes?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

and this... \[\Large \sqrt{(2b - a+b)^2 + (0-c)^2}\] simplifies into\[\Large \sqrt{(3b - a)^2 + (-c)^2}\] But \[\Large (-c)^2=c^2\] so...?

OpenStudy (anonymous):

(-c)^2 = c^2? why? wouldnt it be just c?

terenzreignz (terenzreignz):

Because it isn't? LOL Try -4. \[\Large (-4)^2 =_? \ 4 \qquad \color{red}?\] I don't think so :P \[\Large (-4)^2 = \color{blue}{4^2}\]

OpenStudy (anonymous):

True ._. I feel blonde.

terenzreignz (terenzreignz):

LOL so this ... \[\Large \sqrt{(3b - a)^2 + (-c)^2}\]is just \[\Large \sqrt{(3b-a)^2 +c^2}\] the same as the one earlier, thus showing that they are equal :)

OpenStudy (anonymous):

Okay thank youu!

OpenStudy (anonymous):

Seee, the only problem is if you go and look at the second equation i gave you it is √(2a - a+b)^2 + (0 - c)^2, which would equal √(a+b)^2 + (c)^2?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Yes... what's wrong with that?

OpenStudy (anonymous):

Then they aren't equal D:

terenzreignz (terenzreignz):

Why not?

OpenStudy (anonymous):

√(a+b)^2 + (c)^2 and √3b - a)^2 + (c)^2

terenzreignz (terenzreignz):

wait, why is it (a+b)^2? It is also (3b - a) check it again...

OpenStudy (anonymous):

√(2a - a+b)^2 + (0 - c)^2 ???? (2a - a+b)^2 equals (a+b)^2 right??

terenzreignz (terenzreignz):

Oh? Could you post the original question again? Just so we're clear, and we'll start over :D

OpenStudy (anonymous):

Prove MD = MU: Original Q: MD = √(2b - a+b)^2 + (2c - c)^2 MD = √3b - a)^2 + (c)^2 Original Q: MU = √(2a - a+b)^2 + (0 - c)^2 MU = √(a + b)^2 + (-c)^2 MU = √(a + b)^2 + (c)^2

terenzreignz (terenzreignz):

Well, it can't be proved unless that isn't all of it...

OpenStudy (anonymous):

D:

terenzreignz (terenzreignz):

Why is it MD and MU anyway?

OpenStudy (anonymous):

I have to prove the diagonal is bisecting DU and M is the midpoint so I need the distance from m and d and m and u to see if theyre equal proving that it is being bisected.

terenzreignz (terenzreignz):

May I see the figure?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

"Prove(5-4): MD=MU and MA=MQ"

terenzreignz (terenzreignz):

LOL this helps :)

OpenStudy (anonymous):

.-.

terenzreignz (terenzreignz):

NO, really :D for one thing, you don't really have to prove that the SQUARE ROOTS of the stuff are equal, you didn't need to bother, you just needed the insides of the square roots to be equal

terenzreignz (terenzreignz):

Okay, so let's get MD^2 instead... (this just removes the radical) \[\Large [2b - (b+a)]^2 +(2c-c)^2\]

terenzreignz (terenzreignz):

Do you see your mistake? You forgot to put a+b inside parentheses...

terenzreignz (terenzreignz):

good think I asked you to show me the figure :P

terenzreignz (terenzreignz):

This now simplifies into... \[\Large (b+a)^2 + c^2\]

terenzreignz (terenzreignz):

And everything's okay now, I hope? Let's check MU

OpenStudy (anonymous):

thats the same thing i said tho. wouldnt it be now (b - a)^2? and c^2

terenzreignz (terenzreignz):

Whoops, I meant.. \[\Large (b-a)^2 + c^2\]

terenzreignz (terenzreignz):

Yeah, so that's. \[\Large MD^2\]

terenzreignz (terenzreignz):

Let's check \(MU^2\)

terenzreignz (terenzreignz):

\[\Large [2a -(a+b)]^2+(0-c)^2\]correct?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, this simplifies into... \[\Large (a-b)^2 + (-c)^2\]

terenzreignz (terenzreignz):

But you know... \[\Large (a-b)^2 = (b-a)^2\] and \[\Large (-c)^2 = c^2\] so.......

OpenStudy (anonymous):

So we have: (b - a)^2 +( c)^2 and (a- b)^2 + (c)^2? whihc are equal?

terenzreignz (terenzreignz):

Yes :) so consequently, since they are equal, their square roots are equal :)

terenzreignz (terenzreignz):

Or you could work them while already under the square root, it makes no difference :)

OpenStudy (anonymous):

lol okay, well now i gotta solve MA = √((2a+2b) - (a+b))^2 + (2c - c)^2 = MQ = √(0 - (a+b))^2 + (0 - c)^2 but i can prolly just go do it on my ownn

terenzreignz (terenzreignz):

Well, just ring me if you run into any more problems :) And for goodness' sake, post the question :P Imagine how hard this'd be if it was all in chat o.O

OpenStudy (anonymous):

:PPP

terenzreignz (terenzreignz):

I cannot believe I haven't fanned you yet -.-